Find the measures of the angles in a triangle if the ratio of the measures of the three angles is 13:16:21.

Thanks this is my last question of the day. :]

Found 2 solutions by stanbon, duckness73:
Answer by stanbon(75887)

Find the measures of the angles in a triangle if the ratio of the measures of the three angles is 13:16:21.
----------------------------------
13:16:21 is the same as 13x:16x:21x
----
Equation:
13x + 16x + 21x = 180
50x = 180
x = 18/5 = 3.6
====================
13x = 46.8 degrees
16x = 57.6 degrees
21x = 75.6 degrees
=========================
Cheers,
Stan H.
============

Answer by duckness73(47)

In a case where you have ratios, you add the ratios together to get the total number of "units". In this case, 13 + 16 + 21 = 50 "units". Your first angle (13 "units") represents (13/50) of the total degrees in a triangle (which is 180). So, the first angle is (13/50) * 180 = 46.8 degrees. Your second angle (16 "units") represents (16/50) of the total degrees in a triangle = (16/50) * 180 = 57.6 degrees. Your last angle (21 "units") represents (21/50) of the total degrees = (21/50) * 180 = 75.6 degrees.
Let's see if we get the total right:
46.8 + 57.6 + 75.6 = 180 which is the right number of degrees in a triangle.
So, the angles are 46.8 degrees, 57.6 degrees, and 75.6 degrees.